# Coordinate space representation of a one dimensional $p$-wave pseudopotential

### v1

@@ IV.3 Soliton wavefunction
with $a_{1D}>0$. In the absence of harmonic confinement ($a_{\mathrm{ho}}\to\infty$), wave function (19) coincides with the wave function of the McGuire soliton [27]. Wave function (19) is an exact solution of Hamiltonian (1) in the presence of an additional linear interaction $\sigma|x|$ with $\sigma=-m\omega g/\hbar$. In the limit of strong attraction, $a_{1D}\to+0$, the soliton size is of the order of $\xi\propto a_{1D}/N$, so that the linear potential $\sigma|x|$ is sampled in the vicinity of its zero value, so its presence does not affect much the solution. In other words, the linear potential can be neglected in the $a_{1D}\to+0$ limit and wave function (19) can be used for describing system properties in $a_{1D}\to 0$ limit.
@@ V Results
In this section, we present the numerically computed energy spectrum obtained using both the continuous and discrete representations of the $p$-wave interaction for fermions confined in a harmonic potential. Furthermore, we calculate the eigenvalues of the one-body density matrix for different particle numbers and scattering lengths.
@@ V.1 Energy spectrum
The energy spectrum of the Hamiltonian Eq. (1) for two fermions is known analytically [32, 33]. The relative energy $E_{rel}$ is obtained by solving the equation
@@
where $a_{\mathrm{ho}}=\sqrt{\hbar/(m\omega)}$ is the harmonic oscillator length. Both representations of the $p$-wave interaction introduced above allow us to compute the energy of this system. Although the continuous representation corresponds to a finite interaction range, as this range decreases, the energy must converge to Eq. (21).
@@
Figure 2: Energy spectrum of the relative motion of two $p$-wave fermions in a $1D$ harmonic trap as a function of the interaction parameter. The solid gray line shows the odd solutions in the zero-range model, as given by Eq. (21), while the dotted black line corresponds to the energy obtained by using the discrete representation of the $p$-wave potential (9), the two curves fully coincide. Results for two models of finite-range continuous potentials are shown for potential range equal to $R=0.2,\,0.5,\,1.0\,a_{\mathrm{ho}}$ are shown in blue, red, and green, respectively. Dashed lines correspond to the calculations with the square well and the dash-dotted lines to the modified Pöschl–Teller potentials.
@@
Figure 3: Relative wave functions of two $p$-wave fermions in a $1D$ harmonic trap. Panel (a) shows the three lowest relative wave functions for $a_{1D}=-a_{\mathrm{ho}}$, computed analytically and with the discrete representation. Panel (b) compares the ground state relative wave function for $a_{1D}=-a_{\mathrm{ho}}$ using the discrete and continuous representations.
@@
Figure 2 shows the energy spectrum in the zero-range limit. The analytical solution is shown with a solid gray line, while the numerical results obtained using the discrete representation are shown as a dotted black line. Both results are in complete agreement, demonstrating that the discrete representation, indeed, correctly reproduces the physics of the $p$-wave interaction in the zero-range limit. Let us now examine how the energy depends on the interaction strength, expressed as the ratio of the harmonic oscillator length $a_{\mathrm{ho}}$ and the one-dimensional scattering length $a_{1D}$. In the non-interacting limit ($a_{\mathrm{ho}}/a_{1D}\to+\infty$), the ground-state energy of the relative motion approaches $E_{\text{rel}}=(3/2)\hbar\omega$, corresponding to the occupation of the two lowest single-particle levels with energies $(1/2)\hbar\omega$ and $(3/2)\hbar\omega$, after subtracting the $(1/2)\hbar\omega$ center-of-mass contribution. As the strength of the attractive interaction is increased, the energy decreases monotonically. In the Fermionic Tonks-Girardeau limit, the scattering length diverges and the interaction parameter passes through zero, $a_{\mathrm{ho}}/a_{1D}\to 0$. At that point, the fermionic system has the same energy as two ideal bosons, so that $E_{\text{rel}}=\hbar\omega/2$. For positive values of the scattering length, in the continuum, a bound state enters at the threshold $1/a_{1D}=0$, with energy $E_{b}=-\hbar^{2}/(ma_{1D}^{2})$. In the presence of harmonic confinement, however, no bound state exists at the value of the continuum threshold and the energy is positive. Still, for sufficiently small $a_{1D}>0$, the relative energy tends to the dimer energy $E_{\text{rel}}\to E_{b}$ as the bound state becomes more deeply bound and its reduced spatial extent makes it increasingly insensitive to the external confinement.
@@
For the continuous finite-range interaction potentials, the energy approaches the zero-range limit as the interaction range is reduced, $R\to 0$. In dashed lines are shown the results using the square well potential while in dash-dotted lines are for the modified Pöschl–Teller potential. For a given interaction range, the results obtained for the square well potential are more similar to Eq. (21) than the ones obtained using the modified Pöschl–Teller potential.
@@
The kinetic energy is directly related to the curvature of the wave function with larger curvature corresponding to a higher kinetic energy. The system energy $E_{\text{rel}}$ is increased as additional nodes are introduced in the relative wave function. However, due to its odd parity, each excitation adds two nodes. Consequently, the $n$-th excited relative wave function has $2n+1$ nodes. The ground state for fermions has one node in its relative wave function, the first excited state has three nodes, the second has five, and so on. Figure 3(a) shows the three lowest eigenfunctions for $a_{1D}=-a_{\mathrm{ho}}$ computed both analytically and with the discrete representation. The results confirm this predicted behavior, with the numerically obtained wave functions for the ground and excited states closely matching the analytical solutions. Figure 3(b) compares the ground state wave function for $a_{1D}=-a_{\mathrm{ho}}$ obtained with the continuous and discrete representations. We observe that as the interaction range $R_{SW}$ decreases, the eigenfunctions from the continuous representation progressively converge toward the discrete solution.
@@ V.2 One-body density matrix and natural orbitals
Since the energy spectrum of bosons with $s$-wave interactions is identical to that of fermions with $p$-wave interactions, provided the interaction strengths are appropriately matched, both systems share the same energetic and diagonal properties. However, their off-diagonal properties are significantly different. It is therefore of particular interest to compute the one-body density matrix (OBDM) and perform a natural orbital analysis. The OBDM is defined as
@@
Figure 4(a) shows the largest eigenvalues of the OBDM for the ground state of two fermions as a function of the interaction parameter. All the eigenvalues of the are doubly degenerate (i.e., there exist two orthogonal natural orbitals sharing the same occupation number). In the non-interacting limit, the two largest eigenvalues converge to unity, while the others approach zero. As the inverse of the scattering length goes from -$\infty$ to 0, the two largest eigenvalues decrease and the remaining eigenvalues increase. In the FTG limit $1/a_{1D}\to 0$ the eigenvalues have the analytical form [34, 35]
@@
where $n=1,2,\ldots$. If the inverse of the interaction strength is further increased, the two largest eigenvalues decrease to their minima while the others reach their maxima. In this regime, there is a strong attraction between the particles, so a dimer is formed. The relative wave function has a pronounced peak at $x_{1}=x_{2}^{\pm}$, meaning that knowledge of the position of one particle fully determines the others. We see that the eigenvalues of the OBDM approach the ones found by using the soliton-like wave function (19).
@@
As the inverse of the scattering length further increases, multiple eigenvalues contribute significantly. In this limit, the eigenvalues of Eq. (19) are difficult to compute due to limited numerical resolution. We only show the convergence of the largest eigenvalue of the soliton model.
@@
Figure 4: Eigenvalues of the OBDM for the ground state of a system of NN $p$-wave fermions confined in a harmonic trap. Panels (a) and (b) show the ten (for $N=2$) and nine (for $N=3$) largest eigenvalues, respectively, as a function of the interaction parameter. Numerical results are depicted as solid lines. In dotted lines are shown the results using Eq. (19), while the analytical FTG-limit [34] and non-interacting-limit values are indicated by dashed and dash-dotted lines, respectively. In panel (a), all eigenvalues are doubly degenerate, whereas in panel (b), the largest eigenvalue is non-degenerate and the rest are doubly degenerate. Panel (c) shows the OBDM eigenvalues for N=2, 3, 4, 5N=2,\,3,\,4,\,5 at an intermediate scattering length, $a_{1D}=-a_{\mathrm{ho}}$. Cross markers represent results obtained via exact diagonalization for $p$-wave interacting fermions, using a wave function mapped from the corresponding $s$-wave bosonic system, while circular markers with error bars correspond to Monte Carlo results.
Moreover, for the case of two and three fermions, we validate our numerical results using two independent methods. The first method exploits the mapping between $p$-wave fermions and $s$-wave bosons [23]. The spectrum of $s$-wave bosons with interaction strength $g_{\text{B}}$ in a harmonic trap is equivalent to that of $p$-wave fermions with interaction strength $g_{\text{F}}=-4\hbar^{4}/(m^{2}g_{\text{B}})$. The eigenfunctions of both systems are related as
$\Psi_{\text{F}}(x_{1},\ldots,x_{N})=\mathcal{A}(x_{1},\ldots,x_{N})\Psi_{\text{B}}(x_{1},\ldots,x_{N}),$ (25)
where $\mathcal{A}$ is the unit antisymmetric function
$\mathcal{A}(x_{1},\ldots,x_{N})=\prod_{i<j}^{N}\text{sgn}(x_{i}-x_{j}).$ (26)
@@
For a small number of particles, it is feasible to discretize phase space and perform exact diagonalization. However, since the size of the Hilbert space grows exponentially with the number of particles $N$, this approach becomes impractical for larger systems. To overcome this limitation, we employ the Diffusion Monte Carlo (DMC) method (see Ref. [36] for a general reference). In our case, the fermionic wave function can be mapped to that of bosons, for which the complexity of the DMC algorithm scales as $N^{2}$, allowing us to efficiently access larger particle numbers. The guiding wave function used in Monte Carlo simulations is chosen by applying the antisymmetrization (26) to the bosonic wave function (20). The diffusion Monte Carlo method corrects the wave function and we use extrapolation to calculate the one-body density matrix.
@@
Figure 4(c) reports the seven largest OBDM eigenvalues for N=2, 3, 4, 5N=2,\,3,\,4,\,5 particles for a characteristic value of the repulsive interactions, $a_{1D}=-a_{\mathrm{ho}}$, shown with a dashed line in Figs 4(a,b). For even number of particles, all eigenvalues are doubly degenerate, while for an odd number of particles, all eigenvalues are doubly degenerate except for the largest one, which is non-degenerate. While bosons generally tend to occupy the lowest single-particle state, this is not the case for fermions. For the interaction strength $a_{1D}=-a_{\mathrm{ho}}$, we find that the lowest $n=1,\dots,N$ orbitals are strongly populated, with weights of order unity, whereas the occupation of higher orbitals is strongly suppressed. This behavior is characteristic of weakly interacting fermions. In contrast, in the fermionic Tonks–Girardeau regime ($a_{\mathrm{ho}}/a_{1D}=0$), the occupations follow a qualitatively different distribution, scaling as $\propto 1/n^{2}$ for large $n$ and large particle numbers [34].
@@ VI Conclusions
In this work, we have studied a system of $N$ spin-aligned fermions with $p$-wave interaction confined to a one-dimensional harmonic trap. We have introduced a discrete representation of the $p$-wave interaction, which has the correct zero-range limit and validated it by comparing its results with analytical solutions and with two continuous interaction models based on a square well and a modified Pöschl–Teller potential. Finally, we have performed natural orbital analysis and have calculated the eigenvalues of the one-body density matrix as a function of the one-dimensional scattering length.
Beyond the specific results presented here, these representations provide an alternative method to study systems with $p$-wave interaction. Moreover, they can be used to numerically investigate fermion properties in more complex scenarios, such as in the presence of external fields.

### v2

@@ IV.3 Soliton wave function
with $a_{1D}>0$. In the absence of harmonic confinement ($a_{\mathrm{ho}}\to\infty$), wave function Eq. (20) coincides with the wave function of the McGuire soliton [24]. Wave function Eq. (20) is an exact solution of Hamiltonian Eq. (1) in the presence of an additional linear interaction $\sigma|x|$ with $\sigma=-m\omega g/\hbar$. In the limit of strong attraction, $a_{1D}\to+0$, the soliton size is of the order of $\xi\propto a_{1D}/N$, so that the linear potential $\sigma|x|$ is sampled in the vicinity of its zero value, so its presence does not affect much the solution. In other words, the linear potential can be neglected in the $a_{1D}\to+0$ limit and wave function Eq. (20) can be used for describing system properties in $a_{1D}\to 0$ limit.
@@
The wave functions Eqs. (20-21) are symmetric under exchange of two particles, and, thus, satisfy Bose-Einstein statistics. The fermionic symmetry can be imposed by multiplying the wave functions by the antisymmetric function, $\Psi_{F}=\mathcal{A}\Psi_{B}$, where
$\mathcal{A}(x_{1},\ldots,x_{N})=\prod_{i<j}^{N}\text{sgn}(x_{i}-x_{j}),$ (22)
which corresponds to Girardeau mapping [14]. Defined in this way, the fermionic wave functions $\Psi_{F}$ change the sign, whenever positions of two particles are exchanged.
For a state for which the many-body wave function is known, Monte Carlo methods can be efficiently used to sample its properties, especially, when it can be written in a pair-product for, as in Eqs. (21-22). We will use it for the calculation of the OBDM and its eigenvalues.
@@ V Results
In this section, we present the numerically computed energy spectrum obtained using both the discrete and continuous representations of the interaction for fermions confined in a harmonic potential. Furthermore, we calculate the eigenvalues of the one-body density matrix for different particle numbers and scattering lengths.
@@ V.1 Energy spectrum
The energy spectrum of the Hamiltonian Eq. (1) for two fermions is known analytically [7, 20]. The relative energy $E_{\text{rel}}$ is obtained by solving the equation
@@
where $a_{\mathrm{ho}}=\sqrt{\hbar/(m\omega)}$ is the harmonic oscillator length. Both representations of the interaction introduced above allow us to compute the energy of this system. Although the continuous representation corresponds to a finite interaction range, as this range decreases, the energy must converge to Eq. (23).
@@
Figure 2: Energy spectrum of the relative motion of two fermions in a $1D$ harmonic trap as a function of the interaction parameter. The solid gray line shows the odd solutions in the zero-range model, as given by Eq. (23), while the dotted black line corresponds to the energy obtained using the discrete representation of the potential Eq. (9) with a grid spacing of $\Delta x=0.01\,a_{\mathrm{ho}}$ over the interval $x\in[-10,10]\,a_{\mathrm{ho}}$. The two curves coincide perfectly. Results for two models of finite-range continuous potentials are shown for potential range equal to $R=0.2,\,0.5,\,1.0\,a_{\mathrm{ho}}$ are shown in blue, red, and green, respectively. Dashed lines correspond to the calculations with the square-well and the dash-dotted lines to the modified Pöschl–Teller potentials.
@@
Figure 3: Relative wave functions of two fermions in a $1D$ harmonic trap. Panel (a) shows the three lowest relative wave functions for $a_{1D}=-a_{\mathrm{ho}}$, computed analytically and with the discrete representation. Panel (b) compares the ground state relative wave function for $a_{1D}=-a_{\mathrm{ho}}$ using the discrete and continuous representations.
@@
Figure 2 shows the energy spectrum in the zero-range limit. The analytical solution is shown with a solid gray line, while the numerical results obtained using the discrete representation are shown as a dotted black line. Both results are in complete agreement, demonstrating that the discrete representation, indeed, correctly reproduces the physics of the odd-parity interaction in the zero-range limit. Let us now examine how the energy depends on the interaction strength, expressed as the ratio of the harmonic oscillator length $a_{\mathrm{ho}}$ and the one-dimensional scattering length $a_{1D}$. In the non-interacting limit ($a_{\mathrm{ho}}/a_{1D}\to+\infty$), the ground-state energy of the relative motion approaches $E_{\text{rel}}=(3/2)\hbar\omega$, corresponding to the occupation of the two lowest single-particle levels with energies $(1/2)\hbar\omega$ and $(3/2)\hbar\omega$, after subtracting the $(1/2)\hbar\omega$ CM contribution. As the strength of the attractive interaction is increased, the energy decreases monotonically. In the Fermionic Tonks-Girardeau limit, the scattering length diverges and the interaction parameter passes through zero, $a_{\mathrm{ho}}/a_{1D}\to 0$. At that point, the fermionic system has the same energy as two ideal bosons, so that $E_{\text{rel}}=\hbar\omega/2$. For positive values of the scattering length, in the continuum, a bound state enters at the threshold $1/a_{1D}=0$, with energy $E_{b}=-\hbar^{2}/(ma_{1D}^{2})$. In the presence of harmonic confinement, however, no bound state exists at the value of the continuum threshold and the energy is positive. Still, for sufficiently small $a_{1D}>0$, the relative energy tends to the dimer energy $E_{\text{rel}}\to E_{b}$ as the bound state becomes more deeply bound and its reduced spatial extent makes it increasingly insensitive to the external confinement.
@@
For the continuous finite-range interaction potentials, the energy approaches the zero-range limit as the interaction range is reduced, $R\to 0$. In dashed lines are shown the results using the square-well potential while in dash-dotted lines are for the modified Pöschl–Teller potential. For a given interaction range, the results obtained for the square-well potential are more similar to Eq. (23) than the ones obtained using the modified Pöschl–Teller potential.
@@
The kinetic energy is directly related to the curvature of the wave function with larger curvature corresponding to a higher kinetic energy. The system energy $E_{\text{rel}}$ is increased as additional nodes are introduced in the relative wave function. However, due to its odd parity, each excitation adds two nodes. Consequently, the $n$-th excited relative wave function has $2n+1$ nodes. The ground state for fermions has one node in its relative wave function, the first excited state has three nodes, the second has five, and so on. Figure 3(a) shows the three lowest eigenfunctions for $a_{1D}=-a_{\mathrm{ho}}$ computed both analytically and with the discrete representation. The results confirm this predicted behavior, with the numerically obtained wave functions for the ground and excited states closely matching the analytical solutions. Figure 3(b) compares the ground state wave function for $a_{1D}=-a_{\mathrm{ho}}$ obtained with the continuous and discrete representations. We observe that as the interaction range $R$ decreases, the eigenfunctions from the continuous representation progressively converge toward the discrete solution.
@@ V.2 One-body density matrix and natural orbitals
Since the energy spectrum of bosons with even-parity interactions is identical to that of fermions with odd-parity interactions, provided the interaction strengths are appropriately matched, both systems share the same energetic and diagonal properties. However, their off-diagonal properties are significantly different. It is therefore of particular interest to compute the one-body density matrix (OBDM) and perform a natural orbital analysis. The OBDM is defined as
@@
Figure 4(a) shows the largest eigenvalues of the OBDM for the ground state of two fermions as a function of the interaction parameter. All the eigenvalues are doubly degenerate (i.e., there exist two orthogonal natural orbitals sharing the same occupation number). In the non-interacting limit, the two largest eigenvalues converge to unity, while the others approach zero. As the inverse of the scattering length goes from -$\infty$ to 0, the two largest eigenvalues decrease and the remaining eigenvalues increase. In the FTG limit, $1/a_{1D}\to 0$, the eigenvalues have the analytical form [30, 21]
@@
where $n=1,2,\ldots$. If the inverse of the interaction strength is further increased, the two largest eigenvalues decrease to their minima while the others reach their maxima. In this regime, there is a strong attraction between the particles, so a dimer is formed. The relative wave function has a pronounced peak at $x_{1}=x_{2}^{\pm}$, meaning that knowledge of the position of one particle fully determines the others. We see that the eigenvalues of the OBDM approach the ones found by using the soliton-like wave function Eq. (20) properly antisymmetrized.
@@
As the inverse of the scattering length further increases, multiple eigenvalues contribute significantly. In this limit, the OBDM eigenvalues of Eq. (20) are difficult to compute due to limited numerical resolution. We only show the convergence of the largest eigenvalue of the soliton model.
@@
Figure 4: Eigenvalues of the OBDM for the ground state of a system of $N$ fermions confined in a harmonic trap. Panels (a) and (b) show the ten (for $N=2$) and seven (for $N=3$) largest eigenvalues, respectively, as a function of the interaction parameter. Numerical results using the discrete representation of the odd-parity potential are depicted as solid lines. The results obtained using Eq. (20) are shown with dotted lines. The analytical results for the two limiting cases of zero and infinite scattering length [30] are included for comparison. Specifically, the values in the FTG limit ($a_{ho}/a_{1D}=0$) are represented by crosses, while the asymptotic values in the non-interacting limit ($a_{ho}/a_{1D}\to\infty$) are indicated by arrows. In panel (a), all eigenvalues are doubly degenerate, whereas in panel (b), the largest eigenvalue is non-degenerate and the rest are doubly degenerate. Panel (c) shows the OBDM eigenvalues for N=2, 3, 4, 5N=2,\,3,\,4,\,5 at an intermediate scattering length, $a_{1D}=-a_{\mathrm{ho}}$. Cross markers represent results obtained using the ground state wave function obtained using the discrete representation of the odd-parity potential function Eq. (9) for N=2, 3N=2,\,3, while circular markers with error bars correspond to Monte Carlo results.
@@
Moreover, for the case of two and three fermions, we validate our numerical results using two independent methods. The first method exploits the discrete representation of the potential Eq. (9). For a small number of particles, it is feasible to discretize the coordinate space and diagonalize the Hamiltonian. However, since the size of the position space grows exponentially with the number of particles $N$, this approach becomes impractical for larger systems. To overcome this limitation, we employ the Diffusion Monte Carlo (DMC) method (see Ref. [31] for a general reference). In our case, the fermionic wave function can be mapped to that of bosons, for which the complexity of the DMC algorithm scales as $N^{2}$, allowing us to efficiently access larger particle numbers. The guiding wave function used in Monte Carlo simulations is chosen by applying the antisymmetrization Eq. (22) to the bosonic wave function Eq. (21). The diffusion Monte Carlo method corrects the wave function and we use extrapolation to calculate the one-body density matrix.
@@
Figure 4(c) reports the seven largest OBDM eigenvalues for N=2, 3, 4, 5N=2,\,3,\,4,\,5 particles for a characteristic value of the repulsive interactions, $a_{1D}=-a_{\mathrm{ho}}$, shown with a dashed line in Figs 4(a,b). For an even number of particles, all eigenvalues are doubly degenerate, while for an odd number of particles, all eigenvalues are doubly degenerate except for the largest one, which is non-degenerate. While bosons generally tend to occupy the lowest single-particle state, this is not the case for fermions. For the interaction strength $a_{1D}=-a_{\mathrm{ho}}$, we find that the lowest $n=1,\dots,N$ orbitals are strongly populated, with weights of order unity, whereas the occupation of higher orbitals is strongly suppressed. This behavior is characteristic of weakly interacting fermions. In contrast, in the fermionic Tonks–Girardeau regime ($a_{\mathrm{ho}}/a_{1D}=0$), the occupations follow a qualitatively different distribution, scaling as $\propto 1/n^{2}$ for large $n$ and large particle numbers [30].
@@ VI Conclusions
In this work, we have studied a system of $N$ spin-aligned fermions with odd-parity interaction confined to a one-dimensional harmonic trap. We have introduced a discrete representation of the interaction, which has the correct zero-range limit and validated it by comparing its results with analytical solutions and with two continuous interaction models based on a square-well and a modified Pöschl–Teller potential. Finally, we have performed natural orbital analysis and have calculated the eigenvalues of the one-body density matrix as a function of the one-dimensional scattering length.
Beyond the specific results presented here, this representation provides an alternative method to study systems with odd-parity interaction with the advantage that it eliminates the finite-range effects that arise when using the Pöschl–Teller or square-well potentials. Moreover, these methods can be used to numerically investigate fermion properties in more complex scenarios, such as in the presence of external fields.
@@ VII Acknowledgments
Appendix A Error analysis
Figure 5: Absolute value of the energy difference $\Delta E\equiv|E_{\mathrm{Anl.}}-E_{\mathrm{Num}}|/(\hbar\omega)$ between the zero-range analytical result, Eq. (23), and the numerical results obtained using discrete and continuous representations, as a function of the grid spacing $\Delta x$ with $x\in[-10,10]\,a_{\mathrm{ho}}$. Green squares, blue stars, and red circles correspond to results obtained using the discrete representation and the continuous representation with a square-well and a modified Pöschl–Teller potential, respectively. For the continuous representations, the interaction range is fixed to $R=50\Delta x$. The convergence is analyzed for four scattering lengths: $a_{1D}/a_{\mathrm{ho}}=-5,\,-1,\,1,\,5$, shown in panels (a), (b), (c), and (d), respectively.
In this section, we present an error analysis of the energy spectrum obtained using the discrete and continuous representations of the odd-parity pseudopotential. To this end, we consider the relative Hamiltonian of two interacting fermions in a harmonic trap,
$\hat{\mathcal{H}}_{\text{rel}}=-\frac{\hbar^{2}}{m}\frac{d^{2}}{dx^{2}}+\frac{1}{4}m\omega^{2}x^{2}-g_{\text{F}}\delta^{\prime}(x)\hat{\partial}_{x}.$ (27)
To assess the accuracy of the different representations, we first analyze the convergence of the numerical results using the discrete representation as the grid spacing $\Delta x$ is varied. As shown in Fig. 5, the ground state energy obtained from the numerical simulations approaches the analytical results, Eq. (23), as the grid spacing is reduced. We study the convergence for four different characteristic values of the scattering lengths. By performing a linear fit to the energy, $\Delta E/(\hbar\omega)=a\Delta x/a_{\mathrm{ho}}+b$, we find that in the limit of small grid spacing, $\Delta x\to 0$, the discrepancy is as small as $\Delta E/(\hbar\omega)\sim 10^{-5}$ for all considered cases. This result demonstrates the high accuracy of the discrete representation and its systematic convergence toward the zero-range solution.
While the discrete representation provides a systematic way to converge to the zero-range solution by refining the grid, alternative approaches model the interaction using finite-range continuous potentials. These representations, Eqs. (15,19), depend explicitly on the interaction range, $R$. For a given grid spacing, the range of the interaction cannot be chosen arbitrarily small, since a sufficient number of grid points must lie within the potential region to properly capture the underlying physics. In our calculations, we therefore fix the interaction range to $R=50\Delta x$. As shown in Fig. 5, the energies obtained using these continuous representations also converge to the zero-range solution, Eq. (23) as the grid spacing is reduced, while maintaining the condition $R\gg\Delta x$. However, when comparing the three methods at the same computational cost (i.e., using the same grid spacing), the discrete representation is found to be significantly more accurate than the continuous ones. Specifically, we observe a ratio $\Delta E_{\mathrm{cont}}/\Delta E_{\mathrm{disc}}\sim 10^{2}$, indicating that continuous representations require substantially finer grids to achieve comparable accuracy. This increased computational demand limits the number of particles that can be efficiently simulated using finite-range continuous potentials.
For scattering lengths $a_{1D}/a_{\mathrm{ho}}=-5,\,-1,\,5$, the ratio $\Delta E_{\mathrm{cont}}/\Delta E_{\mathrm{disc}}\sim 10^{2}$, confirming the superior accuracy of the discrete representation in these regimes. In contrast, for $a_{1D}/a_{\mathrm{ho}}=1$, the difference between the discrete and continuous approaches becomes even more pronounced, with $\Delta E_{\mathrm{cont}}/\Delta E_{\mathrm{disc}}\sim 10^{3}$. In this regime, the relative wave function exhibits a large amplitude near the origin, indicating that the relevant physics is dominated by short-range behavior. As a consequence, the finite-range nature of the continuous representations makes it increasingly difficult to accurately capture the ground-state energy, in comparison with the discrete representation. This observation suggests that the discrete approach effectively suppresses finite-range effects inherent to square-well and modified Pöschl–Teller potentials. As the scattering length becomes smaller but remains positive, the wave function amplitude near the origin increases further, exacerbating this issue and leading to larger discrepancies for the continuous representations, as illustrated in Fig. 2.